A brand of soap has a lead time of eight weeks to reorder, a weekly average usage of 20 units and a weekly standard deviation of 8 units. The reorder point for this soap, assuming a 99% service level (z-score = 2.33)
Answer #1

Average weekly demand (d) = 20 units

Standard deviation of weekly demand (\sigmad) = 8 units

Lead time (L) = 8 weeks

At 99% service level value of Z = 2.33

Reorder point = d × L + (Z × \sigma d × √L)

= 20 × 8 + (2.33 × 8 × √8)

= 160 + (2.33 × 8 × 2.83)

= 160 + 52.75

= 212.75 or rounded to 213 units